OpenStudy (anonymous):

Solve the equation. Check your solution. 4 = √h - 6

jhonyy9 (jhonyy9):

squareroot of h or h-6 ?

OpenStudy (anonymous):

100

OpenStudy (anonymous):

10=√h, square both sides.

OpenStudy (anonymous):

so its c?

jhonyy9 (jhonyy9):

ok so like a first step add to the both sides 6 check what will get

OpenStudy (haseeb96):

B is the correct answer see here how? B:-100 4=sqrt h -6 4=sqrt100 -6 4=10-6 4=4